Write down the parametric coordinates of any point on each of the following parabolas.

(a) y2 = 8x

b) y2 = 24x

c) y2 = -16x

d) x2 = 6y


1
Expert's answer
2021-08-15T17:49:48-0400

(a) The parametric coordinates of any point on parabola "y^2 = 8x" :



"\begin{cases}
 x = \dfrac {t^2} 8 \\
 y = t 
\end{cases} \"




"-\infin < t < \infin"



(b) The parametric coordinates of any point on parabola "y^2 = 24x":



"\begin{cases}
 x = \dfrac {t^2} {24} \\
 y = t 
\end{cases} \"




"-\infin < t < \infin"



(c) The parametric coordinates of any point on parabola "y^2 = - 16 x" :



"\begin{cases}
 x = -\dfrac {t^2} {16} \\
 y = t 
\end{cases} \"




"-\infin < t < \infin"



(d) The parametric coordinates of any point on parabola "x^2 = 6y" :



"\begin{cases}
 x = t \\
 y = \dfrac {t^2} 6 
\end{cases} \"




"-\infin < t < \infin"